# Michel F. C. Haddad

### v1

@@ Abstract
We propose a difference-in-differences (DiD) method for a time-varying continuous treatment and multiple time periods. Our framework assesses the average treatment effect on the treated (ATET) when comparing two non-zero treatment doses. The identification is based on a conditional parallel trend assumption imposed on the mean potential outcome under the lower dose, given observed covariates and past treatment histories. We employ kernel-based ATET estimators for repeated cross-sections and panel data adopting the double/debiased machine learning framework to control for covariates and past treatment histories in a data-adaptive manner. We also demonstrate the asymptotic normality of our estimation approach under specific regularity conditions. In a simulation study, we find a compelling finite sample performance of undersmoothed versions of our estimators in setups with several thousand observations.
@@ 1 Introduction
We suggest novel semiparametric kernel-based DiD estimators for both repeated cross-sections, in which subjects differ across time periods; and panel data, in which the same subjects are repeatedly observed across periods. The estimators adopt the double/debiased machine learning (DML) framework introduced in Chernozhukov et al. (2018) to control for covariates and past treatment histories in a data-adaptive manner, which appears attractive in applications where the number of potential control variables is large relative to the sample size. Thus, we estimate the models for the conditional mean outcomes and conditional treatment densities - also known as generalized treatment propensity scores, based on machine learning and use these nuisance parameters as plug-ins in doubly robust (DR) score functions (see for instance Robins et al. (1994) and Robins and Rotnitzky (1995), which are tailored to DiD-based ATET estimation).
@@
These DR scores include kernel functions on the continuous treatment as also considered in Kennedy et al. (2017) and Colangelo and Lee (2023), when imposing a selection-on-observables-assumption (rather than a parallel trend assumption) for identification. Under an appropriately shrinking bandwidth as the sample size increases, we demonstrate that our estimators satisfy Neyman-orthogonality, implying that they are substantially robust to approximation errors in the nuisance parameters under specific regularity conditions - e.g., approximate sparsity when using lasso regression for nuisance parameter estimation as considered in Belloni et al. (2014)). Following Chernozhukov et al. (2018), we also apply cross-fitting such that the nuisance parameters and scores are not simultaneously estimated in the same parts of the data, to avoid overfitting. To account for the biases introduced by the kernel functions, we use undersmoothed bandwidths, implying that the biases vanish asymptotically. We show that under particular regularity conditions, the DML kernel estimators satisfy asymptotic normality, which also provides us with an asymptotic formula for variance estimation. We also conduct a simulation study, which points to a compelling finite sample behavior of undersmoothed versions of our estimators under a sample size of several thousand observations.
@@
Our study adds to a growing literature on the semi- and nonparametric DiD-based identification of causal effects of continuously distributed treatments. This framework avoids potential misspecification errors of classical linear two-way fixed effects models - see, for instance, the discussion in De Chaisemartin and d’Haultfoeuille (2020). Callaway et al. (2024) consider a treatment that has a mass at zero (constituting control group), although might also take continuously distributed non-zero doses. This permits identifying the ATET of a specific non-zero treatment dose versus no treatment, under the parallel trend assumption that the treatment group (receiving a non-zero dose) and the control group share the same trend in mean outcomes under non-treatment. However, Callaway et al. (2024) does not consider semi- or nonparametric estimation conditional on covariates. Zhang (2024) adopts a conditional parallel trends assumption in the continuous treatment context, implying that parallel trends are assumed to hold conditional on observed covariates, as previously introduced by Abadie (2005) for binary treatments. Zhang (2024) studies the canonical two-period model (with one pre- and one post-treatment period) in both the panel and repeated cross-section data, providing identification, estimation, and inference results using the double/debiased machine learning (DML) framework.
@@
Our paper extends these results in several ways. First, we permit comparisons between two non-zero treatment doses under our stronger parallel trends assumption. Second, we allow for more than two time periods with possibly non-zero treatment doses that may vary over time, such as in a staggered adoption scenario - as e.g., considered for binary treatments in Goodman-Bacon (2021), Sant’Anna and Zhao (2020), Sun and Abraham (2021), and Borusyak et al. (2024)), which motivates conditioning on past treatment histories when assessing the ATET. Third, our approach for repeated cross-sections is capable of appropriately tackling covariate distributions that change within treatment groups over time, while many established approaches require them to be time-invariant within treatment groups, as pointed out in Hong (2013), which concerns e.g., the DID methods for binary or discrete treatments in Abadie (2005) using inverse probability weighting, Sant’Anna and Zhao (2020) using DR, or Chang (2020) using DML. A notable exception is the DML approach in Zimmert (2020), which allows for time-varying covariate distributions within treatment groups, nonetheless applying to binary rather than continuous treatments.
@@
de Chaisemartin et al. (2024) consider treatments with strictly non-zero doses, and discuss effect identification and estimation under a parallel trend assumptions related to those considered in the present paper. The authors focus on evaluating an average ATET across the treatment margins of all subjects of changing their treatment dose, which are considered as treatment dose - e.g., what is the average effect of marginally increasing the vaccination rate on mortality among all regions that increase the vaccination rate by possibly different margins. Conversely, our study suggests estimators for assessing the ATET based on comparing two specific treatment doses rather than averaging across many doses - e.g., the effect of a vaccination rate of 40% versus 30%. As a further distinction, de Chaisemartin et al. (2024) do not consider controlling for covariates when imposing parallel trends, while our estimators permit controlling for covariates in a data-driven manner through machine learning.
The remainder of this paper is organized as follows. Section 2 discusses the DiD-based identification of the ATET with a continuous treatment for repeated cross-sectional data. Section 3 considers identification in panel data. Section 4 proposes estimators based on DML with cross-fitting. Section 5 provides asymptotic results for the proposed estimators. Section 6 presents a simulation study for the repeated cross-section and panel data cases. Section 7 concludes.
2 Identification in repeated cross-sections
@@
This section discusses DiD-based identification in repeated cross-sectional data. Let us first consider a continuously distributed treatment, denoted by $D$, with probability mass at $D=0$. In the first case, we assume that there is a single point in time at which the treatment switches from zero to a positive treatment intensity for a subset of the population (i.e., the treated subjects), while it remains at zero for another subset which constitutes the control group. Accordingly, let $T$ denote the time period, where $T=0$ is the baseline period prior to treatment assignment, while $T>0$ is a post-treatment period after the introduction of a positive treatment dose among the treated. The treatment dose may differ across subjects, although in a first step it is assumed to remain constant across post-treatment periods. $Y_{t}$ denotes the observed outcome in period $T=t$, while $Y_{t}(d)$ is the potential outcome in period $t$ when hypothetically setting the treatment to $D=d$, see e.g., Neyman (1923) and Rubin (1974) for further discussion on the potential outcome framework. Throughout the present work, we impose the stable unit treatment value assumption (SUTVA), meaning that the potential outcomes of a subject are not affected by the treatment of others - see, for instance, the discussion in Rubin (1980) and Cox (1958). This permits defining the average treatment effect on the treated (ATET) of treatment dose $d$ in period $t$ as
$\Delta_{d,t}=E[Y_{t}(d)-Y_{t}(0)|D=d].$ (1)
@@
Furthermore, we denote by $X$ observed covariates that may serve as control variables. We impose the following identifying assumptions, which are closely related to those in Lechner (2011), although referring to a continuous (instead of binary) treatment and multiple post-treatment periods.
@@
(Conditional parallel trends under non-treatment):
$E[Y_{t}(0)-Y_{0}(0)|D=d,X]=E[Y_{t}(0)-Y_{0}(0)|D=0,X]$
for all $d>0$ in the support of $D$ and $t>0$ in the support of $T$.
@@
$\Pr(D=d,T=t|X,(D,T)\in\{(d^{\prime},t^{\prime}),(d,t)\})<1$
for all $d>0$ in the support of $D$, $t>0$ in the support of $T$, and $(d^{\prime},t^{\prime})\in\{(d,0),(0,t),(0,0)\}$.
@@
Assumption 1 imposes parallel trends in the mean potential outcome under non-treatment across treated groups, receiving a non-zero treatment dose in the one hand and the control group with $D=0$ on the other hand, conditional on covariates $X$. Assumption 2 imposes that conditional on $X$, the ATET must be zero in the baseline period, which on average rules out non-zero anticipation effects among the treated in period $T=0$ regarding the treatment to be applied. Assumption 3 requires that for any treated subject in a post-treatment period $t>0$, there exist in terms of $X$ comparable treated subjects in the baseline period, non-treated subjects in the post-treatment period, and non-treated subjects in the baseline period. Lastly, Assumption 4 imposes that the covariates are not causally affected by the treatment. We note that may be a relevant restriction if they include post-treatment variables - as it is frequently the case in repeated cross-sections (see, for instance, the discussion in Caetano et al. (2022).
@@
$E[Y_{t}(d)-Y_{t}(0)|D=d,X]$
$=$ $E[Y_{t}(d)|D=d,X]-E[Y_{0}(0)|D=d,X]-E[Y_{t}(0)|D=d,X]+E[Y_{0}(0)|D=d,X]$
$=$ $E[Y_{t}(d)|D=d,X]-E[Y_{0}(d)|D=d,X]-\{E[Y_{t}(0)|D=d,X]-E[Y_{0}(0)|D=d,X]\}$
$=$ $E[Y_{t}(d)|D=d,X]-E[Y_{0}(d)|D=d,X]-\{E[Y_{t}(0)|D=0,X]-E[Y_{0}(0)|D=0,X]\}$
$=$ $E[Y_{t}|D=d,X]-E[Y_{0}|D=d,X]-\{E[Y_{t}|D=0,X]-E[Y_{0}|D=0,X]\}.$
The first equality in (2) follows from subtracting and adding $E[Y_{0}(0)|D=d,X]$, and the second from Assumption 2, implying that $E[Y_{0}(0)|D=d,X]=E[Y_{0}(d)|D=d,X]$. The third equality follows from Assumption 1. The fourth equality follows from the fact that $Y_{t}=Y_{t}(d)$ conditional on $D=d$, which is known as the observational rule. By Assumption 4, the effect of the treatment on the outcome $Y_{t}$ is not mediated by $X$, such that the conditional ATET in expression (2) captures the total causal effect among the treated given $X$, rather than only a partial effect not mediated by $X$. Averaging the conditional ATET over the distribution of $X$ among the treated in a post-treatment period $T=t$ yields the ATET in that period:
@@
where we use the short hand notation $\mu_{d}(t,x)=E[Y_{t}|D=d,X=x]$ for the conditional mean outcome given the treatment, the time period, and the covariates.
If the treatment $d$ consisted of discrete treatment doses, the ATET could alternatively be assessed based on the following doubly robust (DR) expression, which is in analogy to Zimmert (2020) for binary treatments:
@@
where $\Pi=\Pr(D=d,T=t)$ denotes the unconditional probability of receiving treatment dose $d>0$, and being observed in period $t>0$. $\rho_{d,t}(X)=\Pr(D=d,T=t|X)$ for $d\geq 0$ and $t\geq 0$ are the conditional probabilities (also known as propensity scores) of specific treatment group-period-combination $D=d,T=t$ given $X$. Lastly, $I\{\cdot\}$ denotes the indicator function, which is equal to one if its argument is satisfied and zero otherwise.
To see that the DR expression (2) corresponds to the ATET in equation (3), note that the first term on the right of equation (2), $E\left[\frac{I\{D=d\}\cdot I\{T=t\}\cdot[Y_{T}-\mu_{d}(0,X)-\mu_{0}(t,X)+\mu_{0}(0,X)]}{\Pi_{d,t}}\right]$, is by the law of itereated expectations equal to $E[\mu_{d}(t,X)-\mu_{d}(0,X)-(\mu_{0}(t,X)-\mu_{0}(0,X))|D=d,T=t]$ and thus, corresponding to the ATET:
@@
Furthermore, by the law of iterated expectations and basic probability theory, we have for any $d^{\prime}$ and $t^{\prime}$ in the support of $D$ and $T$, respectively, that
@@
Consequently, the remaining terms on the right of equation (3) are equal to zero and expressions (2), such that equation (3) are equivalent and both yield the ATET under out identifying assumptions.
To adapt the previous results based on discrete treatments into a continuous treatment, the indicator functions for non-zero treatment doses $d$ need to be replaced by kernel weighting functions with a bandwidth approaching zero to obtain identification. Thus, we denote by $\omega(D;d,h)$ a weighting function that depends on the distance between $D$ and the non-zero reference value $d$ as well as a non-negative bandwidth parameter $h$: $\omega(D;d,h)=\frac{1}{h}\mathcal{K}\left(\frac{D-d}{h}\right)$, with $\mathcal{K}$ denoting a well-behaved kernel function. The closer $h$ is to zero, the less weight is given to larger discrepancies between $D$ and $d$. Furthermore, we note that the previous (conditional) probabilities $\Pi_{d,t}$ and $\rho_{d,t}(X)$ for $d>0$ and $t\geq 0$ correspond to (conditional) density functions if the treatment is continuous. Considering, for instance, the discussion in Fan et al. (1996), the conditional treatment density, also referred to as generalized propensity score in Imai and van Dyk (2004) and Hirano and Imbens (2005), corresponds to $\rho_{d,t}(X)=\lim_{h\rightarrow 0}E[\omega(D;d,h)\cdot I\{T=t\}|X]$. Moreover, $\Pi_{d,t}=\lim_{h\rightarrow 0}E[\omega(D;d,h)\cdot I\{T=t\}]$. In contrast, $\Pi=\Pr(D=0,T=t)$ and $\rho_{0,t}(X)$ are probabilities as adopted in equation (2), due to the fact that $D$ has a mass point at zero. Applying these modifications to DR expression (2) yields
@@
where $\epsilon_{d,t,x}=(Y_{T}-\mu_{d}(t,x))$ denotes a regression residual. It is worth mentioning that equation (3.9) in Zhang (2024) also provides an DR expression for $\Delta_{d,t}$ with continuous treatment doses $d$. However, in contrast to expression (2), the results in Zhang (2024) are not applicable to setups in which the distribution of covariates $X$ changes within treatment groups over time.
In the next step, we extend our framework to allow for time-varying treatment doses, denoting the treatment dose in period $t$ by $D_{t}$. In addition, we denote by $\mathbf{D_{t}}=\{D_{0},...,D_{t}\}$ the sequence of treatment doses up to and including period $t$. We replace Assumption 1, which imposes parallel trends on the mean potential outcome under non-treatment, with Assumption 5, which imposes parallel trends in mean potential outcomes under positive treatment doses.

### v2

@@ Abstract
We propose a difference-in-differences (DiD) framework designed for time-varying continuous treatments across multiple periods. Specifically, we estimate the average treatment effect on the treated (ATET) by comparing distinct non-zero treatment intensities. Identification rests on a conditional parallel trends assumption that accounts for observed covariates and past treatment histories. Our approach allows for lagged treatment effects and, in repeated cross-sectional settings, accommodates compositional changes in covariates. We develop kernel-based ATET estimators for both repeated cross-sections and panel data, leveraging the double/debiased machine learning framework to handle potentially high-dimensional covariates and histories. We establish the asymptotic properties of our estimators under mild regularity conditions and demonstrate via simulations that their undersmoothed versions perform well in finite samples. As an empirical illustration, we apply our estimator to assess the effect of the second-dose COVID-19 vaccination rate in Brazil and find that higher vaccination rates reduce COVID-19-related mortality after a lag of several weeks.
@@ 1 Introduction
We suggest novel semiparametric kernel-based DiD estimators for both repeated cross-sections, in which subjects differ across time periods, and panel data, in which the same subjects are repeatedly observed across periods. The estimators adopt the double/debiased machine learning (DML) framework introduced in Chernozhukov et al. (2018) to control for covariates and past treatment histories in a data-adaptive manner, which appears attractive in applications where the number of potential control variables is large relative to the sample size. Thus, we estimate the models for the conditional mean outcomes and conditional treatment densities - also known as generalized treatment propensity scores, based on machine learning and use these nuisance parameters as plug-ins in doubly robust (DR) score functions (see for instance Robins et al. (1994) and Robins and Rotnitzky (1995), which are tailored to DiD-based ATET estimation).
@@
These DR scores include kernel functions on the continuous treatment as also considered in Kennedy et al. (2017) and Colangelo and Lee (2023), when imposing a selection-on-observables-assumption (rather than a parallel trend assumption) for identification. Under an appropriately shrinking bandwidth as the sample size increases, we demonstrate that our estimators satisfy Neyman-orthogonality, implying that they are substantially robust to approximation errors in the nuisance parameters under specific regularity conditions - e.g., approximate sparsity when using lasso regression (Tibshirani, 1996) for nuisance parameter estimation as considered in Belloni et al. (2014)). Following Chernozhukov et al. (2018), we also apply cross-fitting such that the nuisance parameters and scores are not simultaneously estimated in the same parts of the data, to avoid overfitting. To account for the biases introduced by the kernel functions, we use undersmoothed bandwidths, implying that the biases vanish asymptotically. We show that under particular regularity conditions, the DML kernel estimators satisfy asymptotic normality, which also provides us with an asymptotic formula for variance estimation.
We also conduct a simulation study, which points to a compelling finite sample behavior of undersmoothed versions of our estimators under a sample size of several thousand observations. As an empirical application, we use Brazilian municipality-level panel data spanning March 2020 to May 2022 to estimate the effect of the continuously distributed second-dose COVID-19 vaccination rate on mortality. Comparing vaccination rates of 60% versus 40%, we find that higher vaccination coverage does not have a statistically significant effect on COVID-19-related mortality in the first two weeks. However, it leads to a statistically significant reduction in mortality after approximately four weeks, consistent with the expected time lag between infection and death.
@@
Our study adds to a growing literature on the semi- and nonparametric DiD-based identification of causal effects of continuously distributed treatments. This framework avoids potential misspecification errors of classical linear two-way fixed effects models - see, for instance, the discussion in De Chaisemartin and d’Haultfoeuille (2020). Callaway et al. (2024) consider a treatment that has a mass at zero (constituting control group), but might also take continuously distributed non-zero doses. This permits identifying the ATET of a specific non-zero treatment dose versus no treatment, under the parallel trend assumption that the treatment group (receiving a non-zero dose) and the control group share the same trend in mean outcomes under non-treatment. However, Callaway et al. (2024) do not consider semi- or nonparametric estimation conditional on covariates. Zhang (2025) adopts a conditional parallel trends assumption in the continuous treatment context, implying that parallel trends are assumed to hold conditional on observed covariates, as previously introduced by Abadie (2005) for binary treatments. Zhang (2025) studies the canonical two-period model (with one pre- and one post-treatment period) in both the panel and repeated cross-section data, providing identification, estimation, and inference results using the double/debiased machine learning (DML) framework. In a related work, Hettinger et al. (2025) combine two parallel trends assumptions to identify the average dose effect on the entire treated group in the panel setting. They adopt a similar framework as Kennedy et al. (2017) to establish a multiply-robust estimator.
@@
Our paper extends these results in several ways. First, we permit comparisons between two non-zero treatment doses under our stronger parallel trends assumption. Second, we allow for more than two time periods with possibly non-zero treatment doses that may vary over time, such as in a staggered adoption scenario - as e.g., considered for binary treatments in Goodman-Bacon (2021), Sant’Anna and Zhao (2020), Sun and Abraham (2021), and Borusyak et al. (2024)), which motivates conditioning on past treatment histories when assessing the ATET. Third, our approach for repeated cross-sections is capable of appropriately tackling covariate distributions that change within treatment groups over time, while many established approaches require them to be time-invariant, as pointed out in Hong (2013), which concerns e.g., the DID methods for binary or discrete treatments in Abadie (2005) using inverse probability weighting, Sant’Anna and Zhao (2020) using DR, or Chang (2020) using DML. A notable exception is the DML approach in Zimmert (2020), which allows for time-varying covariate distributions within treatment groups, nonetheless applying to binary rather than continuous treatments.
@@
We also distinguish our results from de Chaisemartin et al. (2024), which considers treatments with strictly non-zero doses, and discuss effect identification and estimation under parallel trend assumptions related to those considered in the present paper. The authors focus on evaluating an average ATET across the treatment margins of all subjects of changing their treatment dose, which are considered as treatment dose - e.g., what is the average effect of marginally increasing the vaccination rate on mortality among all regions that increase the vaccination rate by possibly different margins. In contrast, our study suggests estimators for assessing the ATET based on comparing two specific treatment doses rather than averaging across many doses - e.g., the effect of a vaccination rate of 40% versus 30%. As a further distinction, de Chaisemartin et al. (2024) do not consider controlling for covariates when imposing parallel trends, while our estimators permit controlling for covariates in a data-driven manner through machine learning.
The remainder of the paper is organized as follows. Section 2 motivates and establishes the identification of the instantaneous ATET for continuous treatments in repeated cross-sections. Section 3 generalizes this framework to time-varying treatments, accommodating lagged effects and dynamic responses. Section 4 addresses identification in panel data settings. Section 5 introduces our proposed estimators based on double machine learning (DML) with cross-fitting, while Section 6 derives their asymptotic properties. Section 7 assesses finite-sample performance through a simulation study. Section 8 presents an empirical application evaluating the effectiveness of the second-dose COVID-19 vaccination rate in Brazil on mortality. Section 9 concludes.
2 Instantaneous effects in repeated cross-sections
@@
This section discusses the identification of the instantaneous ATET, evaluated shortly after treatment introduction, in the repeated cross-sectional setting. For intuition, we first consider a discrete treatment, denoted by $D$. We first assume that there is a single point in time at which the treatment switches from zero to a non-zero treatment dose for a subset of the population (i.e., the treated subjects), while it remains at zero for another subset which constitutes the control group. Accordingly, let $T$ denote the time period, where $T=0$ is the baseline period prior to treatment assignment, while $T>0$ is a post-treatment period after the introduction of a positive treatment dose among the treated. The treatment dose may differ across subjects and for now it is assumed to remain constant across post-treatment periods. Let $Y_{t}$ denote the observed outcome at period $T=t$, while $Y_{t}(d)$ is the potential outcome at period $t$ when hypothetically setting the treatment to $D=d$, see e.g., Neyman (1923) and Rubin (1974) for further discussion on the potential outcome framework. Throughout the present work, we impose the stable unit treatment value assumption (SUTVA), meaning that the potential outcomes of a subject are not affected by the treatment of others - see, for instance, the discussion in Rubin (1980) and Cox (1958). This permits defining the average treatment effect on the treated (ATET) of treatment dose $d$ in period $t$ as
$\Delta_{d,t}=E[Y_{t}(d)-Y_{t}(0)|D=d,T=t].$ (1)
@@
Furthermore, let $X$ denote the observed covariates. We impose the following identifying assumptions, which are closely related to those in Lechner (2011).
@@
(Conditional parallel trends):
$E[Y_{t}(0)-Y_{0}(0)|D=d,T=t,X]=E[Y_{t}(0)-Y_{0}(0)|D=0,T=t,X]$
for $d$ in the support of $D$ and $t>0$ in the support of $T$.
@@
[…]
@@
Assumption 1 imposes parallel trends in the mean potential outcome under non-treatment across treated and control groups, conditional on covariates $X$. Assumption 2 imposes that conditional on $X$, the ATET must be zero in the pre-treatment period $T=0$, which on average rules out non-zero anticipation effects among the treated in period $T=0$. Assumption 3 requires that for any subject receiving treatment dose $d$ in a post-treatment period $t>0$, there exist in terms of $X$ comparable treated subjects in the baseline period, non-treated subjects in the post-treatment period, and non-treated subjects in the baseline period. Lastly, Assumption 4 imposes that the covariates are not causally affected by the treatment. We note that this may be a relevant restriction if the researchers include post-treatment variables - as it is frequently the case in repeated cross-sections (see, for instance, the discussion in Caetano et al. (2022).
@@
$E[Y_{t}(d)-Y_{t}(0)|D=d,T=t,X]$
$=$ $E[Y_{t}(d)|D=d,T=t,X]-E[Y_{0}(0)|D=d,T=0,X]$
- $E[Y_{t}(0)|D=d,T=t,X]+E[Y_{0}(0)|D=d,T=0,X]$
$=$ $E[Y_{t}(d)|D=d,T=t,X]-E[Y_{0}(d)|D=d,T=0,X]$
- $\{E[Y_{t}(0)|D=d,T=t,X]-E[Y_{0}(0)|D=d,T=0,X]\}$
$=$ $E[Y_{t}(d)|D=d,T=t,X]-E[Y_{0}(d)|D=d,T=0,X]$
- $\{E[Y_{t}(0)|D=0,T=t,X]-E[Y_{0}(0)|D=0,T=0,X]\}$
$=$ $E[Y_{T}|D=d,T=t,X]-E[Y_{T}|D=d,T=0,X]$
- $\{E[Y_{T}|D=0,T=t,X]-E[Y_{T}|D=0,T=0,X]\}.$
The first equality in (2) follows from subtracting and adding $E[Y_{0}(0)|D=d,X]$, and the second from Assumption 2, implying that $E[Y_{0}(0)|D=d,T-0,X]=E[Y_{0}(d)|D=d,T=0,X]$. The third equality follows from the conditional parallel trends assumption (Assumption 1). The fourth equality follows from the fact that $Y_{T}=Y_{t}(d)$ conditional on $D=d$ and $T=t$, which is known as the observational rule. By Assumption 4, the effect of the treatment on the outcome $Y_{t}$ is not mediated by $X$, such that the conditional ATET in expression (2) captures the total causal effect among the treated given $X$, rather than only a partial effect not mediated by $X$. Averaging the conditional ATET over the distribution of $X$ among the treated in a post-treatment period $T=t$ yields the ATET in that period:
@@
where we use the short hand notation $\mu_{d}(t,x)=E[Y_{T}|D=d,T=t,X=x]$ for the conditional mean outcome given the treatment, the time period, and the covariates.
If the treatment $d$ consisted of discrete treatment doses, the ATET can alternatively be assessed based on the following doubly robust (DR) expression, which is in analogy to Zimmert (2020) for binary treatments:
@@
where $\Pi=\Pr(D=d,T=t)$ denotes the unconditional probability of receiving treatment dose $d$ at period $t>0$, $\rho_{d,t}(X)=\Pr(D=d,T=t|X)$ denotes the conditional probability of a specific treatment group-period-combination $D=d,T=t$ given $X$, and $I\{\cdot\}$ denotes the indicator function.
To see that the DR expression (2) indeed corresponds to the ATET in equation (3), note that the first term on the right of equation (2), by the law of iterated expectation, corresponds to the ATET:
@@
Moreover, again by the law of iterated expectation, we have for any $d^{\prime}$ and $t^{\prime}$ in the support of $D$ and $T$,
@@
Therefore, the remaining terms in equation (2) are all equal to zero. This demonstrates that (3) and (2) are equivalent and both yield the ATET under our identifying assumptions.
To adapt the previous results based on discrete treatments into a continuous treatment, the indicator functions for non-zero treatment doses $d$ need to be replaced by smooth weighting functions with a bandwidth approaching zero to obtain identification. Thus, we denote by $\omega(D;d,h)$ a weighting function that depends on the distance between $D$ and the non-zero reference value $d$ as well as a non-negative bandwidth parameter $h$: $\omega(D;d,h)=\frac{1}{h}\mathcal{K}\left(\frac{D-d}{h}\right)$, with $\mathcal{K}$ denoting a well-behaved kernel function. The closer $h$ is to zero, the less weight is given to realizations of $D$ that are further away from $d$. Moreover, we note that the previous (conditional) probabilities $\Pi_{d,t}$ and $\rho_{d,t}(X)$ for $d>0$ and $t\geq 0$ should be replaced by (conditional) density functions if the treatment is continuous. Note that the conditional treatment density is also referred to as generalized propensity score in the literature, e.g., Imai and van Dyk (2004) and Hirano and Imbens (2005). In fact, as observed in Fan et al. (1996) the conditional densities can also be expressed as the limit of the conditional expectations of kernel functions: $\rho_{d,t}(X)=\lim_{h\rightarrow 0}E[\omega(D;d,h)\cdot I\{T=t\}|X]$. Similarly, $\Pi_{d,t}=\lim_{h\rightarrow 0}E[\omega(D;d,h)\cdot I\{T=t\}]$. Applying these modifications to DR expression (2) yields
@@
where $\epsilon_{d,t,x}=(Y_{T}-\mu_{d}(t,x))$ denotes a regression residual. It is worth mentioning that in a related study, Zhang (2025) also provides an DR expression for $\Delta_{d,t}$ with continuous treatment doses $d$ that in the repeated cross-sections setting. However, in contrast to our results, his approach is not applicable to settings in which the distribution of covariates $X$ changes over time.
Section 3 extends the analysis to the identification of ATETs in later periods following treatment, allowing for time lags and potentially time-varying treatment doses.
3 Time-varying treatment doses in repeated cross-sections
In this section, we extend our DiD framework to accommodate time-varying treatment doses and multiple treatment (and, later, outcome) periods. For notation, we denote the treatment dose in period $t$ by $D_{t}$, and we use $\mathbf{D}_{t}=\{D_{0},...,D_{t}\}$ to denote the sequence of treatment doses up to and including period $t$. We replace Assumption 1, which imposes parallel trends on the mean potential outcome under non-treatment, with Assumption 5, which imposes parallel trends in mean potential outcomes between non-zero treatment doses.